OpenStudy (anonymous):

B is between A and C and AB=8x-3 and BC = 4x+2,and AC=23. find 'x' and AB.

OpenStudy (welshfella):

form an equation 8x - 3 + 4x + 2 = 23 solve for x

OpenStudy (anonymous):

\[8x-3+4x+2=23\]

OpenStudy (welshfella):

can you solve that?

OpenStudy (anonymous):

I know this isn't my question, but I had a similar one. So will someone please tell me if I did this problem correctly? 8x -3 +4x +2 =23 +3 +3 8x +4x +2 =26 -2 -2 8x +4x =24 12x =24 (12x/12) =(24/12) x =2

OpenStudy (welshfella):

yes